OpenStudy (anonymous):

Does the rule y=-3(4^x) represent a function? a) Yes b) No

OpenStudy (anonymous):

I'll give a medal if you help

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks, can you help me with one more

OpenStudy (anonymous):

I't s a function of x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

go on

OpenStudy (anonymous):

ok hold on I have to leave real quick and copy and paste, but just wait

OpenStudy (anonymous):

A 3,300.00 principal earns 4% interest compound annually. After 3 years what is the balance in the account?

OpenStudy (anonymous):

a) 3,712.05 b) 211,200.00 c) 3,696.00 d) 10,296.00

OpenStudy (anonymous):

f(x)=330000*1.04^x where x is the amount of years.

OpenStudy (anonymous):

f(3)=330000*1.04^3=371205

OpenStudy (anonymous):

A right, I though that's what it was, but wasn't sure